# Quandles as pre-Lie skew braces, set-theoretic Hopf algebras &\&& universal ${\cal R}$-matrices

### v1

@@ 5.2. The set-theoretic Hopf algebra
Coassociativity for the above twisted coproducts does not hold any more (even if $(X,\bullet)$ is a group), that is $(\hat{\cal A},\Delta_{F},\epsilon,S)$ is not a Hopf algebra, but it is rather a quasi Hopf-bialgebra (see also [17, 18] for a more detailed discussion on the issue). The identification of the universal associator of the quasi-Hopf algebra will be presented elsewhere.
@@
Let ${\cal P}=\sum_{a,b\in X}e_{a,b}\otimes e_{b,a}$ be the permutation (flip) operator, then the solution of the braid equation is the familiar set-theoretic solution, $r^{F}={\cal P}R^{F}=\sum_{a,b\in X}e_{a,\sigma_{a}(b)}\otimes e_{b,\tau_{b}(a)}.$ In general, from the universal ${\cal R}$-matrix and the Yang-Baxter equation, and after recalling the representations (5.25) and setting $f_{b,a}:=h_{b}w_{a}^{-1},$ $g_{a,b}:=h_{a}q_{\sigma_{a}(b)}w_{\sigma_{a}(b)}$:
$(\pi\otimes\mbox{id}){\cal R}^{F}:=L=\sum_{a,b\in X}e_{b,\sigma_{a}(b)}\otimes g_{a,b},$ $\leavevmode\nobreak\ (\mbox{id}\otimes\pi){\cal R}^{F}:=\hat{L}=\sum_{a,b\in X}f_{b,a}\otimes e_{a,\tau_{b}(a)},$ and
$(\pi\otimes\pi){\cal R}^{F}:=R^{F}=\sum_{a,b\in X}e_{b,\sigma_{a}(b)}\otimes e_{a,\tau_{b}(a)},$ the consistent algebraic relations (5.5)-(5.7) are satisfied (the interested reader is also referred to [18] for detailed computations). These lead to the set-theoretic Yang-Baxter algebra and provide a consistency check on the associated algebraic relations.
@@
[…]
@@
We also recall that $a\circ b=\sigma_{a}(b)\circ\tau_{b}(a)$ for all $a,b\in X.$
@@
    Let $(X,+,\circ)$ be a skew brace and $\sigma_{a}(b)=-a+a\circ b.$ In this case $W_{a^{-1}}=W^{-1}_{a},$ $W_{1}=id_{V}$ and $W_{a}W_{b}=W_{a\circ b}=W_{\sigma_{a}(b)}W_{\tau_{b}(a)},$ i.e. $W$ is a strong $(X,\circ)-$homomorhism. Notice also that if $X$ is two-sided brace, $W_{a}=W_{b}\Rightarrow\sigma_{a}(c)=\sigma_{b}(c)\Rightarrow(a-b)\circ c^{-1}=a-b+c^{-1},$ $\forall c\in X,$ i.e $a-b\in Socle(X).$
@@
    Let $(X,+,\circ)$ be a skew brace and $\sigma^{z}_{a}(b)=-a\circ z+a\circ b\circ z.$ In this case $W_{a}^{T}=W^{-1}_{a}$ (${}^{T}$ denotes transposition), $\sum_{a\in X}e_{a,a}=id_{V}$ and $W_{a}W_{b}=G_{a\circ b}:=\sum_{c\in X}e_{\sigma^{z\circ z}_{a\circ b}(c),c}=W_{\sigma_{a}(b)}W_{\tau_{b}(a)},$ i.e. $W$ is a weak $(X,\circ)-$homomorhism. Notice also that if $X$ is a two-sided brace, $W_{a}=W_{b}\Rightarrow\sigma_{a}(c)=\sigma_{b}(c)\Rightarrow(a-b)\circ c=a-b+c,$ $\forall c\in X,$ i.e $a-b\in Socle(X).$

### v3

@@ 5.2. The set-theoretic algebras
Note that even if $(X,\bullet)$ is a group coassociativity for the above twisted coproducts does not hold any more, that is $(\hat{\cal A},\Delta_{F},\epsilon,S)$ is not a Hopf algebra, but it is rather a quasi Hopf-bialgebra (see also [22, 23] for a more detailed discussion on the issue). The identification of the universal associator of the quasi-Hopf algebra will be presented elsewhere.
@@
Let ${\cal P}=\sum_{a,b\in X}e_{a,b}\otimes e_{b,a}$ be the permutation (flip) operator, then the solution of the braid equation is the familiar set-theoretic solution, $r^{F}={\cal P}R^{F}=\sum_{a,b\in X}e_{a,\sigma_{a}(b)}\otimes e_{b,\tau_{b}(a)}.$ In general, from the universal ${\cal R}$-matrix and the Yang-Baxter equation, and after recalling the representations (5.23) and setting ${\mathrm{f}}_{b,a}:=h_{b}w_{a}^{-1},$ ${\mathrm{g}}_{a,b}:=h_{a}q_{\sigma_{a}(b)}w_{\sigma_{a}(b)}$:
$(\rho\otimes\mbox{id}){\cal R}^{F}:=L^{F}=\sum_{a,b\in X}e_{b,\sigma_{a}(b)}\otimes{\mathrm{g}}_{a,b},$ $\leavevmode\nobreak\ (\mbox{id}\otimes\rho){\cal R}^{F}:=\hat{L}^{F}=\sum_{a,b\in X}{\mathrm{f}}_{b,a}\otimes e_{a,\tau_{b}(a)},$ and
$(\rho\otimes\rho){\cal R}^{F}:=R^{F}=\sum_{a,b\in X}e_{b,\sigma_{a}(b)}\otimes e_{a,\tau_{b}(a)},$ the consistent algebraic relations (5.5)-(5.7) are satisfied (the interested reader is also referred to [23] for detailed computations). These lead to the set-theoretic Yang-Baxter algebra and provide a consistency check on the associated algebraic relations.
@@
We notice for the fundamental representation $\rho:\hat{\cal A}\to\mbox{End}(V),$ $q_{a}\mapsto{\mathrm{q}}_{a}:=\sum_{b\in X}e_{b,a\triangleright b},$ $h_{a}\mapsto e_{a,a},$ $w_{a}\mapsto W_{a}:=\sum_{b\in x}e_{\sigma_{a}(b),b}$
@@
    Let $(X,+,\circ)$ be a skew brace and $\sigma_{a}(b)=-a+a\circ b.$ In this case $W_{a^{-1}}=W^{-1}_{a},$ $W_{1}=id_{V}$ and $W_{a}W_{b}=W_{a\circ b}=W_{\sigma_{a}(b)}W_{\tau_{b}(a)},$ i.e. $W$ is a strong $(X,\circ)-$homomorhism. Notice also that if $X$ is two-sided brace, $W_{a}=W_{b}\Rightarrow\sigma_{a}(c)=\sigma_{b}(c)\Rightarrow(a-b)\circ c^{-1}=a-b+c^{-1},$ $\forall c\in X,$ i.e. $a-b\in Socle(X).$ We also note that if $\tau_{b}(a):=\sigma_{\sigma_{a}(b)}^{-1}(\sigma_{a}(b)\triangleright a)$ and $a\bullet b=a+b,$ $a\triangleright b=-a+b+a,$ then $\tau_{b}(a)=(\sigma_{a}(b))^{-1}\circ a\circ b.$
@@
    Let $(X,+,\circ)$ be a skew brace and $\sigma^{z}_{a}(b)=-a\circ z+a\circ b\circ z.$ In this case $W_{a}^{T}=W^{-1}_{a}$ (T denotes transposition), $\sum_{a\in X}e_{a,a}=id_{V}$ and $W_{a}W_{b}=G_{a\circ b}:=\sum_{c\in X}e_{\sigma^{z\circ z}_{a\circ b}(c),c}=W_{\sigma_{a}(b)}W_{\tau_{b}(a)},$ i.e. $W$ is a weak $(X,\circ)-$homomorphism. Notice also that if $X$ is a two-sided brace, $W_{a}=W_{b}\Rightarrow\sigma_{a}(c)=\sigma_{b}(c)\Rightarrow(a-b)\circ c=a-b+c,$ $\forall c\in X,$ i.e $a-b\in Socle(X).$